Cud you help me solve this question
An urn contains 9 blue, 7 white and 4 black balls. If 2 balls are drawn at random, then what is the probability that only one ball is white?
Many thanks...

Answer by stanbon(75887)

Cud you help me solve this question
An urn contains 9 blue, 7 white and 4 black balls. If 2 balls are drawn at random, then what is the probability that only one ball is white?
----
One Method::
# of ways to succeed: 7*13 = 91
# of random pairs: 20C2 = 190
P(one white) = 91/190
----
A Second Method::
P(2 white) = 7C2/20C2 = 21/190
P(2non-white) = 13C2/20C2 = 78/190
P(one white) = 1 - [(21+78)/20C2] = 1 - 99/190 = 91/190
-----------------------
Cheers,
Stan H.